OpenStudy (rizags):

An 8.5x11-inch sheet of paper is cut in half lengthwise, while an identical sheet of paper is cut in half widthwise. Do the resulting half sheets have the same area, the same perimeter, or both?

OpenStudy (anonymous):

same area, because none of the paper has gone away

OpenStudy (rizags):

when cut in half lengthwise: dimensions become 8.5 x 5.5 When cut in half widthwise: dimensions become 4.25 x 11 correct?

OpenStudy (anonymous):

i know that much

OpenStudy (rizags):

so lengthwise peremeter is 8.5(2)+5.5(2) = 17+11=28 widthwise is 4.25(2)+11(2)=8.5+22=30.5 area is 8.5 x 5.5 = 46.75 area of other one is 4.25 x 11 = 46.75 Is this all correct?